OpenStudy (anonymous):

The volume of a cone is 300 cubic kilometers. If the radius of the base is 9 kilometers, what is the height of the cone? Use 3.14 for π. Show your work and round to the nearest tenth.

OpenStudy (campbell_st):

the volume of a cone is \[V = \frac{1}{3} \pi r^2h\] in your question you have V = 300, pi = 3.14 and r = 9 so you have \[300 = \frac{1}{3} \times 3.14 \times 9^2 \times h\] just solve for h

OpenStudy (anonymous):

Thank you! :)